Find a series solution, centered about to = 0, for the given ODE (1 – 2?)y - 2xy + 2y = 0
Answer #1

(1) Que - Find a series solution, centered about H0 =0, for the given ODE (1-2²) y - 2ay+2y=0 (1-x2)y – 297 + 2y =0 Sulb(3) n [ Ant2 (n+2)(n+1) + n π [ ann (14) no N22 -Q[ annan tae ananzo nal nzo Second series are zero for both n=o and na! andat nzo (3) 92= (6 +0=2) 1.2 9o a -ao at nel az (1+1=2) 2 au o.Cyc 2.3 at naz au = 4+2-2 Q2 = 7 2 12 = 392= 120 3.4 at n=3 مه(4) from (2) - До+ а24 -+ 4-5--+039° +04 2 -- аэ°40, в. — У= Qo 4 42 – а, 2*+ о -- По +o - Torts - Дар“, Со 20 st) = do+ 4 2